A satellite dish in the shape of paraboloid is 10 ft. across, and 4 ft. deep at its vertex. How far is the receiver from the vertex, if it is placed at the focus? Round off your answer to 2 decimal places
1
Expert's answer
2020-11-17T06:25:41-0500

Let c denote the distance between the focus and the vertex.

The equation for a vertical parabola:

"y = \frac{1}{4c}(x-h)^2+k"

Considering the parabola with center at (0, 0):

"y = \frac{1}{4c}x^2"

Now it is given that the dish is 10ft across and 4ft deep at its vertex, this gives a point on the parabola (10, 4):

"4 = \frac{1}{4c}\cdot10^2""\implies" "c=\frac{100}{16}=6.25"


Answer: 6.25ft.